OpenStudy (anonymous):

The measure of a central angle of a regular polygon is 18 degrees, how many sides does it have? a. 20 b. 18. c. 16

jimthompson5910 (jim_thompson5910):

The general rule is Central angle = 360/n where n is the number of sides

OpenStudy (anonymous):

so it's 20?

jimthompson5910 (jim_thompson5910):

You got it. Nice work.

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Can you help with the next one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

What is the measure of each angle in a 12 sided polygon. a. 90 b. 30 c. 150

jimthompson5910 (jim_thompson5910):

In this case, you can use the formula A = 180(n-2)/n where A is the measure of each angle and n is the number of sides. So in this case, n = 12.

OpenStudy (anonymous):

So 150?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Yay! thanks!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

What is an apothem?

jimthompson5910 (jim_thompson5910):

it's the line segment from the center of a polygon to the midpoint of one of the sides

OpenStudy (anonymous):

Find the apothem of a regular polygon with an area of 625m^2 and perimeter of 100 m

jimthompson5910 (jim_thompson5910):

the green line shown in the pic (see the link) is the apothem http://upload.wikimedia.org/wikipedia/commons/thumb/7/73/Apothem_of_hexagon.svg/220px-Apothem_of_hexagon.svg.png

OpenStudy (anonymous):

a. 12 b. 10 c. 12.5

jimthompson5910 (jim_thompson5910):

how many sides does this polygon have? does it say?

jimthompson5910 (jim_thompson5910):

oh wait, nvm I think we can figure that out, one sec

OpenStudy (anonymous):

No. just says it's a regular polygon.

jimthompson5910 (jim_thompson5910):

Well if there are n sides and the perimeter is 100, then each side is 100/n

jimthompson5910 (jim_thompson5910):

Take half that to get (1/2)*(100/n) = 100/(2n) So half of one side is 100/(2n)

jimthompson5910 (jim_thompson5910):

Now if we break up the polygon, we get a bunch of triangles. We get n of them in fact. Here's one such triangle